the length of a rectangular floor is 1 meter less than three times its width if a diagonal of the rectangle is 37 meters find the length of the floor
Answer by jorel1380(3719)

Let w be the width of the floor. Then its length is 3w-1. So:
w²+(3w-1)²=37²
w²+9w²-6w+1=1369
10w²-6w-1368=0
5w²-3w-684=0
(5w+57)(w-12)=0
w=-57/5 or 12
Using the positive value of 12m for the width, we get the length to be 35m. ☺☺☺☺